A graduate student is studying bacteria known to have a growth rate of 5% per day. If the population of a bacterial sample is currently 63,000 bacteria, then how many bacteria will there be in 2 days?


1
Expert's answer
2021-05-07T09:32:12-0400

Solution

Growth rate is 5%, sample size 63000

Using the formula A=P(1+r/n)^nt where A= final amount, P represents the current sample size, r represents the growth rate, n represents the number of times the growth applied per period and t represents the number of time periods.


Thus: A=63000(1+5/100)^2

=63000(1+0.05)^2

=63000(1.05)^2

=69457.5

Since there aren't half bacteria the answer is rounded off to 69458 bacteria

Answer is 69458 bacteria